OpenStudy (anonymous):

f(x)= sqrt(x+2) ; g(x) = 8x - 6 Find f(g(x)).

Parth (parthkohli):

\[f(g(x)) \implies f(8x - 6) \implies \sqrt{8x - 6 + 2} \]

Parth (parthkohli):

Does that help?

OpenStudy (anonymous):

You're awesome

OpenStudy (anonymous):

Thanks so much!

Parth (parthkohli):

You can simplify that a little more.

OpenStudy (anonymous):

Is \[\sqrt{}8x-4\] the simplest form?

OpenStudy (anonymous):

I'm guessing no...

Parth (parthkohli):

Yeah, that is.

Parth (parthkohli):

You can also factor if needed

OpenStudy (anonymous):

It's none of my answer choices/options though :/

Parth (parthkohli):

What are the choices?

OpenStudy (anonymous):

8 sqrt (x-4) is one option

OpenStudy (anonymous):

2 sqrt (2x + 1) is another

OpenStudy (anonymous):

2 sqrt (2x - 1) is another

Parth (parthkohli):

\[\sqrt{8x - 4} \implies \sqrt{2 \times 2(2x - 1)} \implies \cdots\]

OpenStudy (anonymous):

\[f(x)=2\sqrt{2x-1}\] Is that the answer then?

Parth (parthkohli):

Yes!

OpenStudy (anonymous):

Yay!

OpenStudy (anonymous):

Thanks again :)

Parth (parthkohli):

Except that is \(f(g(x)) \)

Parth (parthkohli):

You're welcome again

Parth (parthkohli):

:)

OpenStudy (anonymous):

Oh true haha

OpenStudy (anonymous):

:)